OpenStudy (anonymous):

quadratic equation x^2-3x-18=0

OpenStudy (aroub):

(x-6)(x+3)

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

x=-3, 6

OpenStudy (anonymous):

x^2-6x+3x-18=0 x(x-6)+3(x-6)=0 (x+3)(x-6)=0 x=-3 and x=6